OpenStudy (anonymous):

1H-3 b) How do you get y+1 = Ay - A --> y = (A+1)/(A-1)

OpenStudy (anonymous):

\[y+1 = Ay - A\] re-arranging terms: \[A+1 =Ay-y\] factoring out y from the RHS: \[A+1=y(A-1)\] divide both sides by (A-1): \[\frac{ A+1 }{ A-1 }=y\]

OpenStudy (anonymous):

thanks :)